The ratio of Taylor's CD to Dave's CD is 7 to 8. If they have a total of 150 CD's. How many CD's does Taylor have?
Answer by mananth(16946)

The ratio of Taylor's CD to Dave's CD is 7 to 8. If they have a total of 150 CD's. How many CD's does Taylor have?
Ratio is 7/8
They share a total of 15 parts of 150
1 part = 150/15 = 10
Taylor has 7 parts of the total He has 7*10=70 CD's
AND Dave has 8 partsof the total He has 8*10 = 80 CD's